# Incentive-Aligned Multi-Source LLM Summaries

### v1

@@ 1 Introduction
An off-the-shelf LLM-based summarizer—unable to verify recency or resist instruction-following traps—may end up recommending the amusement park, producing advice that is unsafe.
@@
This form of strategic manipulation is already emerging. Gibney (2025) document preprints that use hidden prompts to steer AI-assisted peer review. Nestaas et al. (2024); Greshake et al. (2023) show that similar tactics apply to LLM-powered search and plugin ecosystems—where carefully crafted website content or plugin docs can boost an attacker’s visibility and even embed instructions in retrieved pages that steer LLM-integrated applications. Together, these findings underscore the need for incentive-robust designs: even when manipulation is possible, it should not be profitable.
Instead of relying on LLM-centric top-layer fixes, we propose an incentive-aligned pipeline that filters sources before summarization (Fig. 1). Our method, Truthful Text Summarization (TTS), works by decomposing documents into atomic claims and using a multi-task peer prediction mechanism (Dasgupta and Ghosh, 2013; Shnayder et al., 2016) to score sources based on informative corroboration. By filtering low-scoring sources, we can generate a summary from a more reliable set of documents, structurally and strategically aligning source incentives with user needs.
Beyond instantiating multi-task peer prediction in the LLM search setting, our formulation adopts several changes to the traditional multi-task peer prediction model. (i) Tasks are endogenous: claims are produced from retrieved text, so we prevent sources from shaping their own evaluation via a leave-one-out construction, and restoring exogeneity for the scored source. (ii) Signals are embedded in prose: stances are conveyed through authored documents and extracted by a LLM; we formalize implementability and an equivalence to the standard signal–report model. (iii) No payments: utility derives from exposure in the AI-generated overview rather than monetary transfers, so we design inclusion based on a threshold cutoff for score that delivers the desired incentive properties.
@@
    An incentive-aligned pipeline for source selection. We design a framework that (i) converts free-form documents into claim-level stances using a leave-one-out construction so sources cannot influence the claims on which they are judged, and (ii) adapts multi-task peer prediction (Dasgupta and Ghosh, 2013; Shnayder et al., 2016) to reward informative corroboration across claims while discounting generic overlap. The resulting scores determine inclusion and weighting in the final summary, tying a source’s visibility to corroborated information and honest reporting. Designed for open-web search where monetary payments are impractical, the mechanism achieves incentive alignment through scoring and inclusion rather than transfers.
@@
Research in RAG looks at similar problems, but largely focuses on optimizing summary quality given a fixed set of sources, without modeling source incentives. Common approaches leverage internal LLM knowledge or strengthen generation via prompting, self-critique, debate, or “LLM-as-a-judge” (Asai et al., 2024; Yan et al., 2024; Wang et al., 2024, 2025). In dynamic domains, however, static priors can hallucinate and lag fast-moving events. We instead focus on an incentive-aligned aggregation mechanism grounded in retrieved evidence. Our framework is flexible enough to also treat the LLM’s internal knowledge as a distinct source, allowing it to be scored and filtered just like any external document.
Concurrently, work on LLM-based peer-informed scoring has split into two directions. One line learns a textual scoring rule aligned to a chosen reference label (e.g., an instructor’s grade), fitting to that external signal (Lu et al., 2025); relatedly, Wu and Hartline (2024) scores text against ground-truth instructor reviews via proper scoring rules implemented with LLM oracles. The second line uses an LLM’s token-level likelihoods to compare reports without gold labels—either by predicting a peer’s text or by estimating dependence with peer references (Lu et al., 2024; Xu et al., 2024). By contrast, we target open-web search, where reference labels are unavailable and likelihood-based comparisons across heterogeneous, noisy, and adversarial pages are brittle: we form leave-one-out atomic claims, extract claim-level stances, and score sources by informative peer agreement before re-summarizing. We present a thorough related works section in Appendix A.
@@ 2 A Model for Truthful LLM Summaries
Summarizing documents directly is risky: language models may be misled and amplify manipulative content over information useful to the reader. We address this by reframing the problem: instead of whole documents, we work with atomic claims extracted from the corpus (e.g., “The Louvre is open on Tuesdays").
To evaluate a source, we generate its claim set from all other sources (leave-one-out, LOO). This prevents a source from shaping the criteria by which it is judged and converts free-form text into a structured, claim-based comparison.
Our approach mitigates manipulation by the combination of (i) LOO-defined atomic claims and (ii) a scoring rule that rewards informative (beyond-chance) agreement. The LOO structure neutralizes prose-level attacks by fixing what is scored, and the scoring rule aligns incentives by valuing corroborated stances over raw consensus.
2.1 High-level overview
Our framework operates in two passes. Given a query $q$, a retrieval step returns a finite set of sources $\mathcal{C}$. Let $\mathcal{T}=\{\tau_{1},\dots,\tau_{|\mathcal{C}|}\}$ denote their documents. The algorithm proceeds as follows:
  1. 1.
    Score each source via leave-one-out (LOO): For each source $\tau_{i}\in\mathcal{T}$:
    1. (a)
      Generate claims: Create a claim set by generating a draft summary from all other sources, and decompose it into atomic claims with a pre-specified LLM-based decomposer $D$.
    2. (b)
      Elicit stances: For each decomposed atomic claim, a pre-specified LLM-based extractor $E$ returns the stance (e.g., supports, contradicts, abstain) for source $i$ and all peers $j\in\mathcal{C}\setminus\{i\}$.
    3. (c)
      Calculate score: Compute the reliability score $\widehat{w}_{i}$ for source $i$ based on its pattern of agreement with peers across the claims. See details in Section 3.
  2. 2.
    Filter and re-summarize: Define reliable sources $\mathcal{T}_{\text{reliable}}=\{\tau_{i}\in\mathcal{T}\mid\widehat{w}_{i}\geq t_{\text{src, i}}\}$, where $t_{\text{src, i}}$ is a predefined inclusion threshold. Generate a final summary formed from the reliable sources.
@@ 2.2 Players and the held-out claim set
To evaluate a given source $i$, we first define the claims on which it will be judged. These claims are formed without using $\tau_{i}$: a summarizer $M$ maps other documents $\{\tau_{j}\}_{j\neq i}$ to a draft, which a decomposer $D$ splits into atomic claims. Because $\tau_{i}$ does not enter construction, the held-out set $T_{i}$ is exogenous to $i$. We score $i$ on all claims in $T_{i}$ and write $K:=|T_{i}|$. Informally, $T_{i}$ is the set of claims induced by query $q$ and the peer documents for $i$ (a “task class” for source $i$). Throughout, we analyze a fixed $i$, and all expectations are taken conditional on $T_{i}$.
@@
Each claim $s_{k}\in T_{i}$ has a true state of correctness, which we model as an unobserved, latent variable $\theta_{k}\in\{0,1\}$ ($1=$ correct, $0=$ incorrect). Conditional on $T_{i}$, we assume a homogeneous class prior $\pi_{i}:=\Pr(\theta_{k}=1\mid T_{i})\in(0,1)$ that is constant across claims $k\in T_{i}$.
@@ 2.4 Signal informativeness, effort, and reporting
We first separate information acquisition from reporting. Each source $i$ chooses effort $e_{i}\in\{0,1\}$. Under effort ($e_{i}=1$), for each claim $k\in T_{i}$, $i$ observes a private binary signal $z_{ik}\in\{0,1\}$ about $\theta_{k}$. Consistent with the exchangeability of claims for source $i$ (Sec. 2.3), we model its signal quality with claim-invariant conditional accuracies on $T_{i}$:
@@
Note that so far we used claim-invariant parametrization for the scored source $i$ (covering $\alpha_{i}$, signal accuracies, and the reporting policy) - this is a convenience justified by exogeneity. We note that this is not strictly required: Appendix H provides a heterogeneous variant with similar guarantees under a stronger but still plausible peer-margin assumption.
@@ 3 Theoretical Analysis
Following standard definitions in multi-task peer prediction (Shnayder et al., 2016; Agarwal et al., 2020), a strategy is uninformed if its report distribution does not depend on the private signal (equivalently, $\eta_{i}=0$). A mechanism is: (i) strongly truthful if the truthful profile strictly dominates every other profile; (ii) informed-truthful if truthful weakly dominates all profiles and strictly dominates any profile with uninformed strategy; and (iii) $\varepsilon$-informed truthful if truthful is within $\varepsilon$ expected utility of any profile and strictly better than any uninformed strategy.
@@ 3.1 Scoring rule and its expectation
We adapt the scoring rule used in multi-task peer-prediction (Dasgupta and Ghosh, 2013; Shnayder et al., 2016) to our setting. Throughout this subsection, we fix a source $i$, condition on query $q$ and its realized held-out pool $T_{i}$, and use a single random permutation $\rho^{(i)}$ of $\{1,\dots,K\}$ (shared across all peers when scoring $i$) to select off-task indices. We assume the number of tasks $K\geq 3$.
@@ 3.3 Strong truthfulness against significant deviations
Although the affine rule gives the cleanest theoretical guarantee, it assumes linear scaling across the full score range, while in practice scores are moderate with clear separation but not extreme. This makes linear mapping brittle, and with deterministic decisions likely preferable in deployments, we adopt a hard threshold for our main text and experiments, and defer the affine result to Appendix F.
@@ 3.4 Finite-$K$: $\varepsilon$-informed truthfulness
The key observation is that the utility error bound $\epsilon$ decreases exponentially as the number of claims $K$ increases, which means even a moderately large number of claims is sufficient to make unwanted deviations unprofitable with very high probability. We discuss this scaling, computational complexity, and other practical implementation details in Appendix G.
@@ 4.1 Experimental setup
We evaluate TTS on 300-sample subsets from two standard information-seeking benchmarks that provide both a concise short answer and a comprehensive long-form answer for each query: Natural Questions (NQ) (Kwiatkowski et al., 2019), which pairs Google queries with annotated Wikipedia answers, and ClashEval (Wu et al., 2024), which covers six topical domains (news, names, locations, years, drugs, records). For each query, we use the long-form answer as ground truth to construct a six-document source pool from the reference answer. This pool contains four reliable sources (three high-fidelity paraphrases and one concise summary) and two unreliable sources that presents a wrong answer (one deceptive, presenting plausible but false information; one adversarial, containing prompt-injection text). Source generation uses gemini-2.5-flash (Comanici et al., 2025); details are in Appendix I.
@@ Methods.
All pipeline steps (claim decomposition, stance extraction, summarization) use gemini-2.5-flash-lite. [^4] We chose the lightweight model to prioritize the low latency and efficiency required for search applications, though the mechanism itself is model-agnostic. This also reflects a realistic asymmetry where attackers can expend more effort than a real-time defense. Appendix I.8 provides ablations with other model combinations. We compare our method, TTS, against three single-pass baselines: Initial Summary (a standard LLM summary of all sources), Majority Prompt (a LLM summary prompted to include only majority claims), and Majority Claims, where an initial LLM summary is decomposed into atomic claims and only claims with majority support are used for another round of re-summary. Unless otherwise specified, we use a fixed global inclusion threshold of $t_{\mathrm{src},i}=0.06$. [^5] In practice, $t_{\mathrm{src},i}$ can be set adaptively by query type and domain (e.g., sports, science, entertainment) to improve performance. In our experiments, we keep a fixed global threshold (0.06) to validate the framework; adaptive thresholding is expected to improve performance, but is orthogonal and left to future work. Details and prompts are in Appendix I.
@@
To measure overall correctness, we report Answer Accuracy, where an LLM judge compares the generated summary against the dataset’s concise short answer. For a more granular analysis, we report claim-level Precision and Recall, using the comprehensive long-form gold answer as the reference. We also include ROUGE/BLEU scores to assess fluency in Appendix I.
4.2 Result 1: Robustness against adversarial and untruthful sources

### v2

@@ 1 Introduction
Crucially, this shift changes the nature of manipulation. In traditional search, results are presented as distinct entries, so a manipulative source is contained to its own slot, naturally limiting the scope of its influence. In summarization, however, sources are active components being synthesized into a single narrative. This integration allows a strategic actor to hijack or disproportionately influence the final answer via prompt injections, misreports, or semantic steering. By doing so, they can potentially override other evidence regardless of their retrieval rank and completely derail the output, making the system far more fragile than traditional search. As a result, relying on static ranking signals like popularity is insufficient; the system needs a mechanism to verify that a source’s claims are corroborated by its peers before they are allowed to shape the summary.
@@
An off-the-shelf LLM-based summarizer, unable to verify recency or resist instruction-following traps, may end up recommending the amusement park, producing advice that is unsafe.
@@
This form of strategic manipulation is already emerging. Gibney [2025] document preprints that use hidden prompts to steer AI-assisted peer review. Nestaas et al. [2024], Greshake et al. [2023] show that similar tactics apply to LLM-powered search and plugin ecosystems—where carefully crafted website content or plugin docs can boost an attacker’s visibility and even embed instructions in retrieved pages that steer LLM-integrated applications. Together, these findings underscore the need for incentive-robust designs: even when manipulation is possible, it should not be profitable.
Instead of relying on LLM-centric top-layer fixes, we propose an incentive-aligned pipeline that filters sources before summarization (Fig. 1). Our method, Truthful Text Summarization (TTS), works by decomposing documents into atomic claims and using a multi-task peer prediction mechanism [Dasgupta and Ghosh, 2013, Shnayder et al., 2016] to score sources based on informative corroboration. By filtering low-scoring sources, we can generate a summary from a more reliable set of documents, structurally and strategically aligning source incentives with user needs.
Our adaptation of multi-task peer prediction to LLM search differs from the standard model in three aspects. First, a source cannot strategically game which tasks it is graded on: the claims used to evaluate a source are generated from the other documents, so its own text cannot shape the evaluation set. Second, sources never submit abstract reports; they author texts that implement their strategies, and an LLM turns that prose into a stance on each claim. We show that these extracted stances are strategically equivalent to the usual signal–report strategies in peer prediction. Third, instead of paying sources, we reward them with visibility and attribution in the generated summary: a source is included in the AI-generated overview only if its peer-prediction score exceeds a threshold, so inclusion (and hence exposure) plays the role of the incentive.
@@
    An incentive-aligned pipeline for source selection. We design a framework that (i) converts free-form documents into claim-level stances using a leave-one-out construction so sources cannot influence the claims on which they are judged, and (ii) adapts multi-task peer prediction [Dasgupta and Ghosh, 2013, Shnayder et al., 2016] to reward informative corroboration across claims while discounting generic overlap. The resulting scores determine inclusion and weighting in the final summary, tying a source’s visibility to corroborated information and honest reporting. Designed for open-web search where monetary payments are impractical, the mechanism achieves incentive alignment through scoring and inclusion rather than transfers.
@@
Research in RAG looks at similar problems, but largely focuses on optimizing summary quality given a fixed set of sources, without modeling source incentives. Common approaches leverage internal LLM knowledge or strengthen generation via prompting, self-critique, debate, or “LLM-as-a-judge” [Asai et al., 2024, Yan et al., 2024, Wang et al., 2024a, 2025a]. In dynamic domains, however, static priors can hallucinate and lag fast-moving events. We instead focus on an incentive-aligned aggregation mechanism grounded in retrieved evidence. Our framework is flexible enough to also treat the LLM’s internal knowledge as a distinct source, allowing it to be scored and filtered just like any external document.
Concurrently, work on LLM-based peer-informed scoring has split into two directions. One line learns a textual scoring rule aligned to a chosen reference label (e.g., an instructor’s grade), fitting to that external signal [Lu et al., 2025]; relatedly, Wu and Hartline [2024] scores text against ground-truth instructor reviews via proper scoring rules implemented with LLM oracles. The second line uses an LLM’s token-level likelihoods to compare reports without gold labels, either by predicting a peer’s text or by estimating dependence with peer references [Lu et al., 2024, Xu et al., 2024]. By contrast, we target open-web search, where reference labels are unavailable and likelihood-based comparisons across heterogeneous, noisy, and adversarial pages are brittle: we form leave-one-out atomic claims, extract claim-level stances, and score sources by informative peer agreement before re-summarizing. We present a thorough related works section in Appendix A.
@@ 2 A Model for Truthful LLM Summaries
Before introducing the full formalism, we provide a high-level picture of the model.
2.1 High-level Overview of the theoretical model
We model each retrieved document as being written by a strategic author. An author can either put in effort (for example, checking facts) or skip that effort. If they do put in effort, they learn something informative about whether each potential claim related to the query is true or false. Crucially, they are not forced to faithfully report what they learned: they can tell the truth, exaggerate, omit information, or even lie. In other words, both the decision to exert effort and the way in which the learned information is presented in the text are strategic choices. Authors care about their documents being seen and clicked, and they trade this off against any cost of effort and strategy. If misrepresenting information could generate more exposure at the same or lower cost, they would be willing to do so.
On the other side, the search or LLM system designer wants the summary shown to the user to consist of correct, helpful, well-researched claims. As an information aggregator rather than a producer, these systems naturally relies on a healthy internet ecosystem where sources exert effort and are honest. The goal, therefore, is to create such a mechanism, where sources are naturally incentivizes an author to do careful research and report what they find truthfully.
Our theoretical model is a stylized abstraction of this interaction. Each source chooses (i) whether to exert effort and obtain an informative private signal about each claim, and (ii) a reporting strategy that maps what it has learned into what it says. For evaluation, the mechanism fixes, for each source, a set of atomic claims generated from the other documents, so a source cannot shape the criteria on which it is judged. It then evaluates sources based on how their stances on these claims line up with those of their peers. Sources whose reports exhibit more informative agreement with peers receive higher scores and are more likely to be included in, and influence, the final LLM-generated overview. In the formal model that follows, we make these objects precise and prove that, under our scoring rule, exerting effort and reporting truthfully is a utility-maximizing strategy.
2.2 Framework Overview
Our framework operates in two passes. Given a query $q$, a retrieval step returns a finite set of sources $\mathcal{C}$ with documents $\mathcal{T}=\{\tau_{1},\dots,\tau_{|\mathcal{C}|}\}$. The algorithm proceeds as follows: (1) Score each source via leave-one-out (LOO): For each source $\tau_{i}\in\mathcal{T}$, we generate a draft summary from all other sources, decompose it into atomic claims using a decomposer $D$, and elicit stances (support/contradict/abstain) from all sources using an extractor $E$. We then compute a reliability score $\widehat{w}_{i}$ based on informative peer agreement (see Section 3). (2) Filter and re-summarize: We define reliable sources as those where $\widehat{w}_{i}\geq t_{\text{src},i}$ and generate the final summary exclusively from this set.
@@ 2.3 Players and the held-out claim set
To evaluate a given source $i$, we first define the claims on which it will be judged. These claims are formed without using source $i$’s own document $\tau_{i}$: a summarizer $M$ maps other documents $\{\tau_{j}\}_{j\neq i}$ to a draft, which a decomposer $D$ splits into atomic claims. Because $\tau_{i}$ does not enter construction, the held-out set $T_{i}$ is exogenous to $i$. We score $i$ on all claims in $T_{i}$ and write $K:=|T_{i}|$. Informally, $T_{i}$ is the set of claims induced by query $q$ and the peer documents for $i$ (a “task class” for source $i$). Throughout, we analyze a fixed $i$, and all expectations are taken conditional on $T_{i}$.
We refer to this as a Leave-One-Out (LOO) construction of the held-out claim set $T_{i}$ for source $i$. In Section 3.5, we show that practically, we do not need to create $|\mathcal{C}|$ such sets. Instead it suffices to divide sources $\mathcal{C}$ into two random subsets $A$ and $B$, any item in set $A$ will be evaluated on the claim set formed by synthesizing documents in set $B$, and vice versa. This preserves all theoretical results, and removes the factor of $|\mathcal{C}|$ in computational complexity. To keep notations simple, however, we still adopt the LOO framework for now and write held-out set as $T_{i}$ for better intuition.
@@
Each claim $s_{k}\in T_{i}$ has a true state of correctness, which we model as an unobserved, latent variable $\theta_{k}\in\{0,1\}$ ($1=$ correct, $0=$ incorrect). Conditional on $T_{i}$, we assume a homogeneous prior $\pi_{i}:=\Pr(\theta_{k}=1\mid T_{i})\in(0,1)$ that is constant across claims $k\in T_{i}$.
@@ 2.5 Signal informativeness, effort, and reporting
We first separate information acquisition from reporting. Each source $i$ chooses effort $e_{i}\in\{0,1\}$. Under effort ($e_{i}=1$), for each claim $k\in T_{i}$, $i$ observes a private binary signal $z_{ik}\in\{0,1\}$ about $\theta_{k}$. Consistent with the exchangeability of claims for source $i$ (Sec. 2.4), we model its observed signal quality with claim-invariant conditional accuracies on $T_{i}$:
@@ 3 Theoretical Analysis
Following standard definitions in multi-task peer prediction [Shnayder et al., 2016, Agarwal et al., 2020], a strategy is uninformed if its report distribution does not depend on the private signal (equivalently, $\eta_{i}=0$). A mechanism is: (i) strongly truthful if the truthful profile strictly dominates every other profile; (ii) informed-truthful if truthful weakly dominates all profiles and strictly dominates any profile with uninformed strategy; and (iii) $\varepsilon$-informed truthful if truthful is within $\varepsilon$ expected utility of any profile and strictly better than any uninformed strategy.
@@ 3.1 Scoring rule and its expectation
We adapt the scoring rule used in multi-task peer-prediction [Dasgupta and Ghosh, 2013, Shnayder et al., 2016] to our setting. Throughout this subsection, we fix a source $i$, condition on query $q$ and its realized held-out pool $T_{i}$, and use a single random permutation $\rho^{(i)}$ of $\{1,\dots,K\}$ (shared across all peers when scoring $i$) to select off-task indices. We assume the number of tasks $K\geq 3$.
@@ 3.3 Strong truthfulness against significant deviations
Although the affine rule gives the cleanest theoretical guarantee, eliminating the need for a large-sample limit, it imposes a continuous linear scaling across the full score range. In practice, our mechanism yields clearly separated clusters of reliable and unreliable sources. A hard threshold directly capitalizes on this clear separation and is therefore preferable. Thus, we utilize the hard threshold for our main analysis and experiments, deferring the affine results to Appendix F.
@@ 3.4 Finite-$K$: $\varepsilon$-informed truthfulness
The key observation is that the utility error bound $\epsilon$ decreases exponentially as the number of claims $K$ increases, which means even a moderately large number of claims is sufficient to make unwanted deviations unprofitable with very high probability. We discuss this scaling and other practical implementation details in Appendix G.
3.5 Computational Complexity
Computing the score $\widehat{w}_{i}$ for one source requires averaging over $K$ claims and $|\mathcal{C}|-1$ peers, resulting in a cost of $O(K(|\mathcal{C}|-1))$. Scoring all sources thus takes $O(|\mathcal{C}|K(|\mathcal{C}|-1))$.
However, as mentioned in Section 2.3, we adopt a simple modification that preserves the key LOO property (each source is evaluated on claims constructed without its own document) while dramatically reducing complexity: instead of constructing $T_{i}$ from $\mathcal{C}\setminus\{i\}$ for every $i$, we randomly split the source set $\mathcal{C}$ into two subsets $\mathcal{C}_{A}$ and $\mathcal{C}_{B}$. For $i\in\mathcal{C}_{A}$, we form the held-out claim set once as $T_{A}=D(M(\mathcal{C}_{B}))$ and evaluate all $i\in\mathcal{C}_{A}$ on $T_{A}$; symmetrically, for $i\in\mathcal{C}_{B}$ we use $T_{B}=D(M(\mathcal{C}_{A}))$. In both cases, the claim set used to score a source is formed from documents that exclude that source, so the exogeneity condition required for our incentive guarantees still holds.
Since we preserve the crucial feature that $T_{i}$ is independent with $\tau_{i}$, all theoretical derivations and theorems follow. This significantly improves the overall computational complexity to only $O(K|\mathcal{C}|)$. In practice, practitioners can choose to only run it on select queries rather than all user queries, and couple it with a reputation prior (compatible with our framework and discussed in Appendix D) to further decrease deployment costs.
@@ 4.1 Experimental Setup
We evaluate TTS on two standard information-seeking benchmarks that provide both a concise short answer and a comprehensive long-form answer for each query: Natural Questions (NQ) [Kwiatkowski et al., 2019], which pairs Google queries with annotated Wikipedia answers, and ClashEval [Wu et al., 2024], which covers six topical domains (news, names, locations, years, drugs, records). For each query, we use the long-form answer as ground truth to construct a six-document source pool from the reference answer. This pool contains four reliable sources (high-fidelity paraphrases), one low quality source (correct answer but very concise with little supporting information), and two unreliable sources that presents a wrong answer (one deceptive, presenting plausible but false information; one adversarial, containing prompt-injection text). Source generation uses gemini-2.5-flash [Comanici et al., 2025]; details are in Appendix H. After preprocessing, there are 1299 samples in ClashEval and 1100 samples in the NQ dataset that we use in our main evaluation.
In addition, to assess TTS on purely real-world evidence, we run a small case study on the Open-Domain QA with Conflicting Contexts dataset [Liu et al., 2025], where all sources are naturally occurring web pages; the construction and results of this experiment are reported in Appendix H.5.
@@ Methods.
All pipeline steps (claim decomposition, stance extraction, summarization) use gemini-2.5-flash-lite. [^4] We chose the lightweight model to prioritize the low latency and efficiency required for search applications, though the mechanism itself is model-agnostic. This also reflects a realistic asymmetry where attackers can expend more effort than a real-time defense. Appendix H.9 provides ablations with other model combinations. We compare our method, TTS (using the computational efficiency-optimized A/B group method descired in Sec. 3.5), against three single-pass baselines: Initial Summary (a standard LLM summary of all sources), Majority Prompt (a LLM summary prompted to include only majority claims), and Majority Claims, where an initial LLM summary is decomposed into atomic claims and only claims with majority support are used for another round of re-summary. Unless otherwise specified, we use a fixed global inclusion threshold of $t_{\mathrm{src},i}=0.06$. [^5] In practice, $t_{\mathrm{src},i}$ can be set adaptively by query type and domain (e.g., sports, science, entertainment) to improve performance. In our experiments, we keep a fixed global threshold (0.06) to validate the framework; adaptive thresholding is expected to improve performance, but is orthogonal and left to future work. Details and prompts are in Appendix H.
@@
To measure overall correctness, we report Answer Accuracy, where an LLM judge compares the generated summary against the dataset’s concise short answer. For a more granular analysis, we report claim-level Precision and Recall, using the comprehensive long-form gold answer as the reference. We also include ROUGE/BLEU scores to assess fluency in Appendix H.
4.2 Results 1: Robustness against adversarial and untruthful sources